OpenStudy (anonymous):

multiply. use the pattern (a+b)(a-b)= a squared - b squared 102 * 98

OpenStudy (anonymous):

@ParthKohli gave you answer to your previous question. This one is just the same. \[(100+2)(100-2)=100^{2}-2^{2}\]

OpenStudy (anonymous):

watching at the nummbers 102 and 98 you have to find two numbers a and b such that: a+b=102 and a-b=98 later: (a+b)(a-b) will be same as write 102*98, but \[(a+b)(a-b)=a ^{2}-b ^{2}\] which is more easy to multiply. That's it